OpenStudy (anonymous):

find the limit. lim approaches +infinity at sin^-1((3x)/(5-6x))

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{3x}{5-6x}=\frac{3}{-6}=\frac{-1}{2}\]

myininaya (myininaya):

actually mik-wind was more correct i think f(x)=sin(x) is defined on (-pi,pi) and its range is (-1,1) so f(x)=sin^{-1}(x) has domain (-1,1) and range (-pi,pi)

myininaya (myininaya):

it is \[\frac{-\pi}{6}\]

OpenStudy (anonymous):

can someone explain to me how the answer is -pi/6

myininaya (myininaya):